OpenStudy (anonymous):

Beth has 3500 feet of fencing available to enclose a rectangular field. One side of the field lies along a river, so only three sides require fencing. Express the area A of the rectangle as a function of x, where x is the length of the side parallel to the river. A(x)=

myininaya (myininaya):

3500=2x+y y=3500-2x A=xy=x(3500-2x) or is it 3500=2y+x y=(3500-x)/2 A=xy=x(3500-x)/2

myininaya (myininaya):

oh wait they say x is parallel to the river so its the 2nd one

myininaya (myininaya):

A=x(3500-x)/2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

just to clarify, that a duck in the bottom of my river.

myininaya (myininaya):

lol cute

OpenStudy (anonymous):

i know right? his name is charles.